A particle is moving in a straight line with a velocity given by 5t2, where t is time. Find an expression for the acceleration of the particle.

Grade:12

1 Answers

Nirmal Singh.
10 years ago
v = 5 t^2
differentiating both side w r to t
dv/dt = 5 d/dt(t^2) = 5*2t
we know a = dv/dt = 10t
Thanks & Regards,
Nirmal Singh
Askiitians faculty